### v1

@@ Introduction
The analysis of multivariate data is frequently carried out in science, and the analysis of bivariate data in particular has established itself as an important field, for example in climate simulations where temperature and pressure are jointly investigated. The relationship between the scalar fields can be examined using Jacobi sets [18] as a tool for analysis. This mathematical concept originates from the field of topology and represents a generalization of critical points to multifields. In topological data analysis, Jacobi sets are used to extract Ridge-Valley graph in computer vision and image processing [31], and to track critical points in time [4, 18]. In addition, they are used to compare scalar fields [19] to study the Reeb space [39] and to drive fiber surface extraction [34]. They are not only used in visualization but also in other scientific disciplines [1, 2, 25, 30]. This multifaceted application shows the importance of Jacobi sets in scientific analysis, which is why these structures must be easy to understand.
@@
However, noise can be a major problem that makes analysis difficult for domain experts. For this reason, some work has already been done on the simplification of Jacobi sets [37, 3]. For example, Reeb graphs are created and the Jacobi sets are simplified using the comparison measure $\kappa$.[37]. This involves removing loops or components of the Jacobi set, which occur in particular in the case of noise and numerical errors. However, only the representation of the Jacobi sets is simplified, the underlying data itself is not adjusted, which is unfavorable for further processing. In contrast, smoothing filters adjust the data globally, but the extent to which important structures are lost in the process has not been sufficiently investigated. Important structures are regions in the dataset that lie above a threshold value of a comparison metric and are also visually separated from the surrounding area , see Topological Simplifcation of Jacobi Sets for Piecewise-Linear Bivariate 2D Scalar Fields with Adjustment of the Underlying Data in the upper left part in the brown region of interest. Subdivision algorithms can also simplify Jacobi sets and, in particular, remove zig-zag, but only lead to a better representation of this [27]. However, the visual evaluation indicates that the number of Jacobi set components of the dataset increases. Therefore, we will compare these approaches and our approach to determine how suitable they are for simplifying the Jacobi sets. Our goal is to modify the input functions in such a way that the extracted Jacobi sets are simplified, while the input functions are only minimally modified. Our approach to simplifying Jacobi sets aims to clarify important structures and make them usable for subsequent analyses by changing the underlying data. This is done by iteratively collapsing cells, where all scalar values in a cell are set to the same value, similar to what is shown in the 1D case in Fig. 2. To identify the components to be collapsed, we use the neighborhood graph of the components of the Jacobi set.
@@ 1 Related Work
An overview of topology-based methods in visualization can be found in the survey by Heine et al. [21]. He et al. [20] focuses more on the visualization of multivariate data. In scalar field topology, the essential features of scalar fields can be described by contour trees [8, 10, 42], Reeb graphs [33] or the Morse-Smale complex [16]. FRWith their method, scalar fields can be topologically simplified and the critical points and their relationships can be reduced. Carr et al. [9] describe a method for simplifying the contour tree by suppressing smaller topological features. Bremer et al. [5] create the Morse-Smale complex of a 2D function and simplify the topology by canceling pairs of critical points. Luo et al. [29] simplify critical points based on a point cloud. Edelsbrunner et al. [17] introduce the idea of persistent homology for the topological simplification of a point cloud, which was introduced by Cohen-Steiner et al. [14]. Edelsbrunner et al. [17] introduce the idea of persistent homology for the topological simplification of a point cloud, which extended Cohen-Steiner et al. [14] to persistence diagrams.
In order to apply the methods mentioned to multifield data, an adaptation is necessary. Methods such as Jacobi sets and Reeb graphs cannot extended so easily. In addition, Carlsson et al. [6] showed that a generalization of persistent homology is difficult. Singh et al. [35] use partial clustering of high-dimensional data and introduce the idea of the mapper based on this. The Joint Contour Net by Carr et al. [7] is partly based on this idea, but use joint contour plates and their topological connectivity. Reeb graphs were introduced in the works [13] and [36] parallel for multifields. Chattopadhyay et al. [13] introduce the Jacobi structure for subdividing the Reeb space, which creates a Reeb skeleton that corresponds to the Reeb graph. An algorithm for calculating the Reeb space of a bivariate, piecewise-linear scalar function on a tetrahedral grid is presented by Tierny et al. [38]. Chattopadhyay et al. [12] use the Joint Contour Net to further simplify multivariate data and to simplify the Reeb skeleton.
@@ 2 Background
[…]
@@ 3.1 Jacobi Set Computation
[…]
@@ 3.2 Neighborhood graph
For the neighborhood graph A, all cells are combined into one node if they are not separated by a Jacobi set. This results in a graph with $69$ nodes and $108$ edges for the domain section under consideration. In neighborhood graph B, as shown in Fig. 1(c), cells are merged into one node if they fulfill the conditions of graph A. In addition, two cells are combined into one node if they are connected via a point and both have a negative orientation. The neighborhood graph C, shown in Fig. 1(d), follows the same approach as graph B, but here two cells are combined into one node if they are connected by a point and both have a positive orientation. These adaptations in graphs B and C are based on the goal of converting the neighborhood graph into a neighborhood tree. Such a tree can be simplified by starting to combine nodes at the leaves as long as the range area of these nodes are smaller than a threshold according to a suitable metric. A side effect of this adaptation is the reduction of nodes in the graph. The effects of these adaptations can be seen in the resulting graphs. The neighborhood graph B has $39$ nodes and $38$ edges, while the graph C has $33$ nodes and $32$ edges. The different orientations lead to significantly different graphs.
Finally, we consider the neighborhood graph D. Here, all cells are combined into one node if they are not separated by a Jacobi set. In addition, two cells are connected if they are connected by a point, have the same orientation and this orientation prevails in the neighborhood of the point. The goal of this approach is to minimize the imbalance that arises in the neighborhood graphs B and C and produce a more analyzable neighborhood tree. The resulting graph for this approach is shown in Fig. 1(e) and has $23$ nodes and $22$ edges. This graph shows a compact structure that may be well-suited for further analysis.
@@ 3.3 Metrics to select Jacobi set components
For the reduction of the Jacobi set components a suitable metric has to be chosen. There are geometric and topological metrics, whereby we have chosen a topological metric in this work. Based on the neighborhood graph, we now decide which of the Jacobi set components should be collapsed. Different measures can be used to choose the components:
  • The Jacobi set component’s domain area can be used:
@@
  • The Jacobi set components range area can also be taken into account:
@@
We chose hypervolume, which takes into account both the area of Jacobi set components in the domain and the range. Furthermore, the other two metrics pose problems, especially if the Jacobi set components areas are very different. This could lead to interesting structures being lost. Another possibility would be the automatic determination of this variable depending on the neighborhood. However, this method was not implemented in this work.
@@ 3.4 Collapse of Regions
To collapse a region, we use a greedy algorithm that shrinks the regions starting from the Jacobi set components by collapsing individual cells until the entire region is collapsed. The collapse a cell in the 2D case is a complex problem due to its influence on the neighborhood. In the 1D case, all values at the vertices of a cell can simply be set to the same value in the range, see Fig. 2. This collapses the line cell to a point in the 1D case. Equivalently, in the 2D case, variant $V1$ mapped all values at the vertices of a triangle cell to one value and thus collapse the cell in the range, as shown in Fig. 3(a). Variant $V2$ collapsed the cell to a line in the range, as shown in Fig. 3(a).
@@
The difference between these two variants is that collapsing to a point affects more neighboring cells compared to collapsing to a line. Therefore, we prefer collapsing to a line in this work. Since only one edge of the triangle cells needs to collapse to collapse the entire cell, we decide which of the edges to select depending on the neighborhood. We can therefore divide this into four cases:
@@ 3.5 Method
Motivated by the theoretical approach of Bhatia et al. [3], we have developed a method that aims to change the functions by collapsing a cell to reduce the complexity of the Jacobi sets while retaining their essential structures. The collapse of a cell means that the area that this cell sets up in the range is reduced to the value 00. This approach simplifies the Jacobi sets and thus also reduces the Jacobi set components. The idea of collapsing comes from the 1D case, in which 2 critical points can be reduced by mapping the function values $f(x)$ of the two points of a cell to the same value. In Fig. 2 this is illustrated for the cell $BC$.
Here you can see that this change affects at least one neighboring cell. Therefore, the choice of a suitable value is also crucial to keep the side effects as low as possible, as otherwise not all critical points can be removed. In the 2D case, this concept is more difficult, as the number of affected neighboring cells increases. Therefore, a suitable metric must be used to keep the side effects as low as possible.
@@ 23                   $break$
This neighborhood graph is used in our method to select Jacobi set components. This is done in FindCollapsibleCells(NG,t), which uses the hypervolume from Section 3.3 and an appropriate threshold to identify nodes that are considered irrelevant. The algorithm is cell-based and therefore returns all cells in a list (CL) that belong to the irrelevant nodes in a list.
The algorithm now begins to process the list CL. For each cell, it calculates whether it lies on the border of a Jacobi set component to be collapsed. To do this, CalculateCellNeighborhood(CL,c) returns for a given cell how many neighboring cells are connected to it and are not contained in CL. This number is used to sort cells and skip cells that are not on the border. For the remaining cells, the list (CV) is returned that contains all possibility variants of how the cell can be collapsed. These variant are described in more detail in Section 3.4.
This list is now used to select the best variant for collapsing the cell. For this purpose, a metric is used that decides which variant has the least side effects based on the direct point neighborhood of the cell. This metric counts how many neighboring cells are negatively affected when the different options are applied. A negative influence occurs when cells change their orientation in the range, as shown by the dashed line in Fig. 2 for the 1D case. In addition, the area in the range in the neighborhood is also taken into account, and changes are preferred if the total area in the range becomes smaller.
After selecting the best variant, the bivariate scalar field is adjusted accordingly in ApplyCollapseVariant(F,c,bcv) according to the Section 3.4. The cell is then removed from the list CL. Since the algorithm cannot always select a variant in which no side effects occur, FlippedCellNeighbors(NG,c) recognizes negatively influenced cells and then adds them to the list CL.
@@ 3.6 Jacobi set Visualization
The Jacobi sets are calculated as described in Section 3.1. In addition, degenerate cells are assigned based on their point neighborhood of the Jacobi set components that predominate in the neighborhood. This means that the cell is assigned to the orientation that is larger when the neighboring cell orientations are summed up. If there are only degenerated cells in the direct neighborhood, the neighbors of the degenerated cells are considered recursively until a clear assignment is possible.
@@ 4.2 Cylinder Flow (Synthetic)
When observing the datasets, three regions of interest (ROI) can be identified that are relevant for the comparison. The first ROI is marked green in the datasets and contains the two right corners where many small Jacobi set components can be recognized in Topological Simplifcation of Jacobi Sets for Piecewise-Linear Bivariate 2D Scalar Fields with Adjustment of the Underlying Data of the original data. These are due to noise, as the range area is very small, which can also be seen from the saturation of the colors. In the loop subdivision, the Jacobi set components can still be recognized very well; a simplification of the Jacobi sets is therefore not possible. After smoothing with the binomial filter, a reduction of the Jacobi set components can already be recognized. The smoothing is further intensified with the Gaussian filter, which even leads to the removal of all Jacobi set components in this ROI. A similar result is obtained after applying the CA variants. Here, all Jacobi set components can also be removed, which can be seen in Fig. 3(h), e, f. Only in variant A do individual Jacobi set components remain. The second ROI is marked in brown in the top center of the original data. A larger reddish Jacobi set component can be seen here, with several smaller components adjacent to it , which should be removed. After applying the loop subdivision, the large Jacobi set component is visible and the small ones are no longer recognizable. When smoothing with the binomial filter, there is hardly any effect in Fig. 3(f) and the large and small Jacobi set components are still present. In contrast, the effect of smoothing by the Gaussian filter is too large and all components are removed, which is visible in Fig. 3(g). When using the CA variant D, the result is similar to the binomial filter, and all components are still present. CA variants B and C do better here, where the large Jacobi set component remains and almost all small components disappear. In CA variant A, only the large component remains, as can be seen in Topological Simplifcation of Jacobi Sets for Piecewise-Linear Bivariate 2D Scalar Fields with Adjustment of the Underlying Data at the bottom left. The third ROI is marked in yellow in the center of the original data. In this ROI, several small Jacobi set components can be recognized in the original data. In the loop subdivision, several large Jacobi set components are separated from each other as can be seen in Fig. 3(e). This indicates that the small components are caused by noise or numerical errors and should disappear. With the smoothing filters, the result is similar to before. The binomial filter smoothes too little and the Gaussian filter too much, whereby in this ROI the Gaussian filter does not remove all the small Jacobi set components, but even creates many new ones. Using the CA variants produces very different results here. In variant C, for example, the Jacobi set components are not separated but connected. In variants B and D, the larger components are separated from each other, but many small components remain. In variant A, the large components of the Jacobi set are completely separated and almost all small components disappear so that the results of the loop subdivision are very similar.
@@
Overall, the visual comparison of the cylinder flow shows that the CA variant A is the best variant. This is also shown in the Section 4. To confirm this, the Topological Simplifcation of Jacobi Sets for Piecewise-Linear Bivariate 2D Scalar Fields with Adjustment of the Underlying Data on the right shows the neighborhood graphs of the Jacobi set components for the original data and for the dataset after applying CA variant A. Here we can see that the algorithm greatly simplifies the Jacobi set components. In addition, the range of the original data is compared with the dataset after applying the CA variant A in Fig. 5 to examine the effects of modifying the data. Despite the differences in the domain, there is hardly any visual difference between the two datasets. The simplification from the Jacobi set components from $679$ to $17$ using the CA variant A takes $0.2$ s.

### v2

@@ Introduction
The analysis of multivariate data is frequently carried out in science, and the analysis of bivariate data in particular has established itself as an important field, for example in climate simulations where temperature and pressure are jointly investigated. The relationship between the scalar fields can be examined using Jacobi sets [18] as a tool for analysis. This mathematical concept originates from the field of topology and represents a generalization of critical points to multifields. In topological data analysis, Jacobi sets are used to extract Ridge-Valley graph in computer vision and image processing [31], and to track critical points in time [4, 18]. Besides, they are used to compare scalar fields [19] to study the Reeb space [39] and to drive fiber surface extraction [34]. They are not only used in visualization but also in other scientific disciplines [1, 2, 25, 30]. This multifaceted application shows the importance of Jacobi sets in scientific analysis, which is why these structures must be easy to understand.
@@
However, noise can be a major problem that makes analysis difficult for domain experts. For this reason, a work on the simplification of Jacobi sets [37] has already been carried out, the Reeb graphs created and the Jacobi sets simplified with the comparison measure $\kappa$. This involves removing loops or Jacobi set components, which occur in particular in the case of noise and numerical errors. However, only the representation of the Jacobi sets is simplified, the underlying data itself is not adjusted, which is unfavorable for further processing. In contrast, smoothing filters adjust the data globally, but the extent to which important structures are lost in the process has not been sufficiently investigated. Important structures are regions in the dataset that lie above a threshold value of a comparison metric and are also visually separated from the surrounding area , see Topological Simplifcation of Jacobi Sets for Piecewise-Linear Bivariate 2D Scalar Fields with Adjustment of the Underlying Data in the upper left part in the brown region of interest. Subdivision algorithms can also simplify Jacobi sets and, in particular, remove zig-zag, but only lead to a better representation of this [27]. However, the visual evaluation indicates that the number of Jacobi set components of the dataset increases. Therefore, we will compare these approaches and our approach to determine how suitable they are for simplifying the Jacobi sets.
Motivated by the theoretical approach of Bhatia et al. [3], we have developed a method that aims to change the functions by collapsing a cell to reduce the complexity of the Jacobi sets while retaining their important structures by changing the underlying data. The collapse of a cell means that the area that this cell sets up in the range is reduced to the value 00, which is also the case with degenerated cells. This approach simplifies the Jacobi sets and thus also reduces the Jacobi set components. The idea of collapsing comes from the 1D case, in which 2 critical points can be reduced by mapping the function values $f(x)$ of the two points of a cell to the same value. In Fig. 2 this is illustrated for the cell $BC$. Here you can see that this change affects at least one neighboring cell. Therefore, the choice of a suitable value is also crucial to keep the side effects as low as possible, as otherwise not all critical points can be removed. In the 2D case, this concept is more difficult, as the number of affected neighboring cells increases. Therefore, a suitable metric must be used to keep the side effects as low as possible. To identify the components to be collapsed, we use the neighborhood graph of the Jacobi set components.
@@ 1 Related Work
An overview of topology-based methods in visualization can be found in the survey by Heine et al. [21]. He et al. [20] focuses more on the visualization of multivariate data.
In scalar field topology, the essential features of scalar fields can be described by contour trees [8, 10, 42], Reeb graphs [33] or the Morse-Smale complex [16]. With their method, scalar fields can be topologically simplified and the critical points and their relationships can be reduced. Carr et al. [9] describe a method for simplifying the contour tree by suppressing smaller topological features. Bremer et al. [5] create the Morse-Smale complex of a 2D function and simplify the topology by canceling pairs of critical points. Luo et al. [29] simplify critical points based on a point cloud. Edelsbrunner et al. [17] introduce the idea of persistent homology for the topological simplification of a point cloud, which was introduced by Cohen-Steiner et al. [14]. Edelsbrunner et al. [17] introduce the idea of persistent homology for the topological simplification of a point cloud, which extended Cohen-Steiner et al. [14] to persistence diagrams.
In order to apply the methods mentioned to multifield data, an adaptation is necessary. Methods such as Jacobi sets and Reeb graphs cannot extended so easily. Besides, Carlsson et al. [6] showed that a generalization of persistent homology is difficult. Singh et al. [35] use partial clustering of high-dimensional data and introduce the idea of the mapper based on this. The Joint Contour Net by Carr et al. [7] is partly based on this idea, but use joint contour plates and their topological connectivity. Reeb graphs were introduced in the works [13] and [36] parallel for multifields. Chattopadhyay et al. [13] introduce the Jacobi structure for subdividing the Reeb space, which creates a Reeb skeleton that corresponds to the Reeb graph. An algorithm for calculating the Reeb space of a bivariate, piecewise-linear scalar function on a tetrahedral grid is presented by Tierny et al. [38]. Chattopadhyay et al. [12] use the Joint Contour Net to further simplify multivariate data and to simplify the Reeb skeleton.
@@ 2 Background
Note that the critical points of $f$ and $g$ are trivially part of their Jacobi set. Edelsbrunner and Harer [18] showed that, if $D$ is a 2-manifold, $J(g,f)$ is generically a collection of pairwise disjoint smooth curves, free of any self-intersections. We will call the connected components of a Jacobi set the Jacobi set components. Eq. 1 has many equivalent formulations, for example, Edelsbrunner et al. [19] mention that for two 2D functions, the Jacobi set is the set of all points where the rank of the Jacobian, i.e., the matrix comprised of the function’s gradients, is smaller than 2. They also present a more general criterion: the set of all points $x$ where the local $\kappa$ measure is 00, where $\kappa(x)$ is defined as the length of the wedge product of the functions gradients. They argue that the average of this measure over an area is related to the amount of effort needed to change the topology of the functions in that area.
@@ 3.1 Jacobi Set Computation
Note that, $\nabla\mathbf{f}(\mathbf{x})$ is called the Jacobian of $\mathbf{f}$ in calculus, and is a matrix composed of the gradients of $f$ and $g$. The absolute value of the Jacobian’s determinant would equal the value of $\kappa$ by Edelsbrunner et al. [19], but by not taking the absolute value, we get an alternative method to compute the Jacobi set for piecewise-linear representations of bivariate 2D fields. For piecewise-linear data the function restricted to the interior of each triangle $\sigma$ can be given as:
@@ 3.2 Neighborhood graph
For the neighborhood graph A, all cells are combined into one node if they are not separated by a Jacobi set. This results in a graph with $69$ nodes and $108$ edges for the domain section under consideration. In neighborhood graph B, as shown in Fig. 1(c), cells are merged into one node if they fulfill the conditions of graph A. Besides, two cells are combined into one node if they are connected via a point and both have a negative orientation. The neighborhood graph C, shown in Fig. 1(d), follows the same approach as graph B, but here two cells are combined into one node if they are connected by a point and both have a positive orientation. These adaptations in graphs B and C are based on the goal of converting the neighborhood graph into a neighborhood tree. Such a tree can be simplified by starting to combine nodes at the leaves as long as the range area of these nodes are smaller than a threshold according to a suitable metric. A side effect of this adaptation is the reduction of nodes in the graph. The effects of these adaptations can be seen in the resulting graphs. The neighborhood graph B has $39$ nodes and $38$ edges, while the graph C has $33$ nodes and $32$ edges. The different orientations lead to significantly different graphs.
Finally, we consider the neighborhood graph D. Here, all cells are combined into one node if they are not separated by a Jacobi set. Besides, two cells are connected if they are connected by a point, have the same orientation and the summed orientation of all point neighbors is also the same. The goal of this approach is to minimize the imbalance that arises in the neighborhood graphs B and C and produce a more analyzable neighborhood tree. The resulting graph for this approach is shown in Fig. 1(e) and has $23$ nodes and $22$ edges. This graph shows a compact structure that may be well-suited for further analysis.
@@ 3.3 Metrics to select Jacobi set components
In this study, the goal is to collapse a selection of Jacobi set components to simplify the data. A suitable metric has to be chosen to select a Jacobi set component. There are geometric and topological metrics, and we have chosen a topological metric in this work. Based on the neighborhood graph, we now decide which of the Jacobi set components should be collapsed. Different measures can be used to choose the components:
  • The Jacobi set component’s domain area:
@@
  • The Jacobi set components range area:
@@
Here $i$ is a cell in a region. We chose hypervolume, which takes into account both the area of Jacobi set components in the domain and the range. Furthermore, the other two metrics pose problems, especially if the Jacobi set components areas are very different. This could lead to important structures being lost. Another possibility would be the automatic determination of this variable depending on the neighborhood. However, this method was not implemented in this work.
@@ 3.4 Collapse of Regions
To collapse a region, we use a greedy algorithm that shrinks the regions starting from the cells at the Jacobi set and collapses them first until the entire region is collapsed. The collapse a cell in the 2D case is a complex problem due to its influence on the neighborhood. In the 1D case, all values at the vertices of a cell can simply be set to the same value in the range, see Fig. 2. This collapses the line cell to a point in the 1D case. In the 2D case, Fig. 3(a) shows that $V1$ maps all values at the vertices of a triangle cell to one value and thus collapses the cell in the range. in Fig. 3(a) $V2$ collapsed the cell to a line in the range.
@@
The difference between these two variants is that collapsing to a point affects more neighboring cells compared to collapsing to a line. Therefore, we prefer collapsing to a line in this work. Since only one edge of the triangle cell needs to collapse to collapse the entire cell, we decide which of the edges to select depending on the neighborhood. In this way, side effects can be reduced, and places where several small Jacobi set components will be removed are collapsed from the outside. Since we know at this point which cells should be collapsed, we check whether there are cells in the neighborhood that should also be collapsed, which we can then divide into four cases:
@@ 23                   $break$
This neighborhood graph is used in our method to select Jacobi set components. This is done in FindCollapsibleCells(NG,t), which uses the hypervolume from Section 3.3 and an appropriate threshold to identify nodes that are considered irrelevant. The algorithm is cell-based and therefore returns all cells in a list (CL) that belong to the irrelevant nodes in a list. The algorithm now begins to process the list CL. For each cell, it calculates whether it lies on the border of a Jacobi set component to be collapsed. To do this, CalculateCellNeighborhood(CL,c) returns for a given cell how many neighboring cells are connected to it and are not contained in CL. This number is used to sort cells and skip cells that are not on the border. For the remaining cells, the list (CV) is returned that contains all possibility variants of how the cell can be collapsed. These variant are described in more detail in Section 3.4. This list is now used to select the best variant for collapsing the cell. For this purpose, a metric is used that decides which variant has the least side effects based on the direct point neighborhood of the cell. This metric counts how many neighboring cells are negatively affected when the different options are applied. A negative influence occurs when cells change their orientation in the range, as shown by the dashed line in Fig. 2 for the 1D case. Besides, the area in the range in the neighborhood is also taken into account, and changes are preferred if the total area in the range becomes smaller. After selecting the best variant, the bivariate scalar field in ApplyCollapseVariant(F,c,bcv) is adjusted. This means that the values of the cell vertices are set to the same value according to the cases from Section 3.4. The cell is then removed from the list CL. Since the algorithm cannot always select a variant in which no side effects occur, FlippedCellNeighbors(NG,c) recognizes negatively influenced cells and then adds them to the list CL.
@@ 3.6 Jacobi set Visualization
The Jacobi sets are calculated as described in Section 3.1. Besides, degenerate cells are assigned based on their point neighborhood of the Jacobi set components that predominate in the neighborhood. This means that the cell is assigned to the orientation that is larger when the neighboring cell orientations are summed up. If there are only degenerated cells in the direct neighborhood, the neighbors of the degenerated cells are considered recursively until a clear assignment is possible.
@@ 4.2 Cylinder Flow (Synthetic)
When observing the datasets, three regions of interest (ROI) can be identified that are relevant for the comparison because they are noisier or the structures are complex. Nevertheless, the rest of the dataset is not uninteresting as important structures should not or only minimally be changed. The first ROI is marked green in the datasets and contains the two right corners where many small Jacobi set components can be recognized in Topological Simplifcation of Jacobi Sets for Piecewise-Linear Bivariate 2D Scalar Fields with Adjustment of the Underlying Data of the original data. These are due to noise, as the range area is very small, which can also be seen from the saturation of the colors. In the loop subdivision, the Jacobi set components can still be recognized very well; a simplification of the Jacobi sets is therefore not possible. After smoothing with the binomial filter, a reduction of the Jacobi set components can already be recognized. The smoothing is further intensified with the Gaussian filter, which even leads to the removal of all Jacobi set components in this ROI. A similar result is obtained after applying the CA variants. Here, all Jacobi set components can also be removed, which can be seen in Fig. 3(h), e, f. Only in variant A do individual Jacobi set components remain. The second ROI is marked in brown in the top center of the original data. A larger reddish Jacobi set component can be seen here, with several smaller components adjacent to it , which should be removed. After applying the loop subdivision, the large Jacobi set component is visible and the small ones are no longer recognizable. When smoothing with the binomial filter, there is hardly any effect in Fig. 3(f) and the large and small Jacobi set components are still present. In contrast, the effect of smoothing by the Gaussian filter is too large and all components are removed, which is visible in Fig. 3(g). When using the CA variant D, the result is similar to the binomial filter, and all components are still present. CA variants B and C do better here, where the large Jacobi set component remains and almost all small components disappear. In CA variant A, only the large component remains, as can be seen in Topological Simplifcation of Jacobi Sets for Piecewise-Linear Bivariate 2D Scalar Fields with Adjustment of the Underlying Data at the bottom left. The third ROI is marked in yellow in the center of the original data. In this ROI, several small Jacobi set components can be recognized in the original data. In the loop subdivision, several large Jacobi set components are separated from each other as can be seen in Fig. 3(e). This indicates that the small components are caused by noise or numerical errors and should disappear. With the smoothing filters, the result is similar to before. The binomial filter smoothes too little and the Gaussian filter too much, whereby in this ROI the Gaussian filter does not remove all the small Jacobi set components, but even creates many new ones. Using the CA variants produces very different results here. In variant C, for example, the Jacobi set components are not separated but connected. In variants B and D, the larger components are separated from each other, but many small components remain. In variant A, the large components of the Jacobi set are completely separated and almost all small components disappear so that the results of the loop subdivision are very similar.
@@
Overall, the visual comparison of the cylinder flow shows that the CA variant A is the best variant. This is also shown in the Section 4. To confirm this, the Topological Simplifcation of Jacobi Sets for Piecewise-Linear Bivariate 2D Scalar Fields with Adjustment of the Underlying Data on the right shows the neighborhood graphs of the Jacobi set components for the original data and for the dataset after applying CA variant A. Here we can see that the algorithm greatly simplifies the Jacobi set components. Besides, the range of the original data is compared with the dataset after applying the CA variant A in Fig. 5 to examine the effects of modifying the data. Despite the differences in the domain, there is hardly any visual difference between the two datasets. The simplification from the Jacobi set components from $679$ to $17$ using the CA variant A takes $0.2$ s.